For a signal x(n)=sin(2*pi*n/5) defined for n=0to7, evaluate the Fast Fourier Transform using signal flow graph. (Use decimation in Time Algorithm).                                               
Answer #1

8 (0) 60. 365 -XCI) - 7,622 -10. 312 न x (2) -२.५१-j2. 125 • ६६) - rcl) - (3) 40. 278 - 51.03 न (५) ८ 0.815 - x() - 0.278 + j(423 4s 3 (บวน 100% - ouds o- () Ups - (1) (r) - * (L) xs - (2) 1 (1) 2 Sn () พ. (s) - shา (f) = 0 2. (6) - ciๆ (47) ) - Sinbis aut al wees 0.59 be = ax + Quy way a 0.95 + 30.36 - ai - acum s-2049 bu 2 a, -Ryu = 0.95 -30.36 bs = qs + ay wag om -0.22